OpenStudy (anonymous):

Time And Work: Question 1: A is 50% more efficient than B. C does half of the work done by A and B together. If C can do a piece of work in 20 days, then how much time will A, B and C take to complete the same work? I am looking for some shortcut trick by which it can be done within 1 minute.. Any suggestions??

OpenStudy (anonymous):

This question most likely can not be done less than 1 minute :)

OpenStudy (kropot72):

The information "A is 50% more efficient than B" is not needed to find the solution. C working alone does 1/20 of the work in 1 day. A and B working together do (1/20) * 2 = 1/10 of the work in 1 day. A, B and C working together do 1/10 + 1/20 = 3/20 of the work in 1 day, and therefore take 20/3 = 6 2/3 days to complete the work.

OpenStudy (anonymous):

Oh My God, how couldn't I notice that.... :( Seriously, a big big blunder from my side.. Yes, the very first line is not needed and the solution will take less than 30 seconds even... :) @kropot72, you are great brother.. :)

OpenStudy (kropot72):

You're very welcome :)